# Optimal Monotone Depth-Three Circuit Lower Bounds for Majority

### v1

@@ 1 Introduction
Paturi, Pudlák, and Zane [13] (henceforth called PPZ) showed that Parity requires $\Sigma^{k}_{3}$ circuits of size $\Omega(2^{n/k})$. In fact, this results in an unrestricted depth-3 lower bound of $\Omega(n^{1/4}2^{\sqrt{n}})$. Both of these lower bounds are known to be tight. Paturi, Pudlák, Saks, and Zane [12] (henceforth called PPSZ) extended the techniques of [13] and showed that recognizing the codewords of a good error-correcting code such as BCH requires $\Sigma^{k}_{3}$ circuits of size $2^{cn/k}$ for a universal constant $c>1$. This remains the best known lower bound for computing any explicit function by $\Sigma^{k}_{3}$ circuits. The core technique in these lower bounds is a combinatorial result about the structure of satisfying assignments of $k$-CNF formulas which can be exploited to solve $k$-SAT.
@@
An algorithm solving $k$-SAT is non-trivial if it has running time $2^{(1-\epsilon_{k})n}$ for some $\epsilon>0$ which we call savings. PPZ and PPSZ used their combinatorial arguments to give non-trivial $k$-SAT algorithms. PPSZ currently holds the record as the best known $k$-SAT algorithm with savings $c/k$ for a universal constant $c>1$ [15]. The Super Strong Exponential Time Hypothesis (SSETH) states that such savings, that is $\Theta(1/k)$, are best possible [18]. We know that indeed this is the case for PPSZ [14].
@@
The Majority function has long been held as a natural candidate to beat the depth-3 barrier [7]. Unlike Parity, the tightness of natural circuit constructions for Majority is not known. To determine the depth-3 complexity of Majority, Gurumukhani et al. [5] suggested a path forward by considering $\Sigma^{k}_{3}$ circuits. Their central conjecture is that the size of any $\Sigma_{3}^{k}$ circuit computing Majority is $2^{\Omega(n\log(k)/k)}$. Together with a standard random restriction argument to reduce the bottom fan-in, this conjecture implies an unrestricted depth-3 lower bound of $2^{\Omega(\sqrt{n\log n})}$ matching a known upper bound [10]. They pointed out that determining $\Sigma^{k}_{3}$ complexity of Majority is non-trivial even for $k=3$, and proposed this as a first step towards settling the depth-3 complexity of Majority. Results of Lecomte, Ramakrishnan, and Tan [11] can be interpreted as proving the conjecture under the assumption that each depth-2 subcircuits depends on at most $k$ input bits.
Similar to PPZ and PPSZ, determining the depth-3 complexity is closely tied to $k$-SAT. Gurumukhani et al. [6] showed that proving the conjectured lower bound for $\Sigma^{k}_{3}$ via a particular approach which they called local enumeration, would break SSETH.
@@
Enum($k$, $t$) is defined as follows. For a natural number $k$ and a parameter $t$, given an $n$-variate $k$-CNF $F$ with no satisfying assignment of Hamming weight less than $t(n)$, enumerate all satisfying assignments of Hamming weight exactly $t(n)$. Analogously, we can define a restriction of this problem to monotone formulas which we denote by Enum+($k$, $t$).
@@ 1.1 Our contribution
Our main result is an optimal lower bound on the size of any monotone $\Sigma^{3}_{3}$ circuit computing Majority matching the upper bounds conjectured in the literature (see e.g. [4]).
Theorem 1.2.
$\operatorname{Size}^{k}_{3}(\operatorname{\mathsf{Maj}}_{n})\leq 2^{O(n\log(k)/k)}$. In particular, $\operatorname{Size}^{3}_{3}(\operatorname{\mathsf{Maj}}_{n})\leq(2/6^{1/4})^{n}\simeq 1.277^{n}$. Furthermore, there are monotone circuits that achieve this bound.
We optimally solve Enum+($3$, $n/2$) to derive our lower bound
@@
While we do not claim that our techniques are new, we believe that our approach to prove matching lower bounds for depth-3 circuits holds promise. The first element of our approach is the top-down technique: if a $\Sigma\Pi\Sigma_{3}^{k}$ circuit computes a Boolean function $f:\{0,1\}^{n}\mapsto\{0,1\}$ and $s$ is the number of its depth-2 circuits, one of the depth-2 circuits which is a $k$-CNF must accept at least $|f^{-1}(1)|/s$ members of $f^{-1}(1)$. This brings our focus to the study of extremal $k$-CNFs with respect to $f$. A $k$-CNF $F$ is extremal with respect to $f$ if ${\textsc{Sat}}(F)\subseteq f^{-1}(1)$ and $|{\textsc{Sat}}(F)|$ is the maximum among such $k$-CNF. Such extremal $k$-CNFs can also used for constructing depth-3 $\Sigma\Pi\Sigma_{3}^{k}$ circuits which are (almost) optimal depending on the symmetries of $f$. Thus, the study of $k$-CNF which are extremal with respect to a Boolean function is of fundamental concern.
@@
What do we know of $k$-CNFs which are extremal with respect to the majority function or more generally with respect to Boolean threshold functions? Our intuition (based on extensive provably optimal constructions for small $n$) suggested that a modularity principle might hold. We say that modularity principle holds for a Boolean function if an extremal $k$-CNF $F$ can be expressed as the conjunction of $k$-CNFs $F^{\prime}_{i}$ where $F^{\prime}_{i}$ is supported on a bounded number of variables and the supports of $F^{\prime}_{i}$ are disjoint. Although we lack a formal method to determine if the majority function is inherently modular, we used the modularity principle to guide a trial-and-error search for constructing putative extremal 3-CNFs of Boolean threshold functions (for thresholds less or equal to $n/2$).
Our final step is to apply induction to prove the optimality of our constructions. Several ingredients helped the induction to succeed: 1. Correct guess (or construction) of optimal constructions for Boolean threshold function with threshold less or equal to $n/2$, 2. Construction of (optimal or near optimal) extremal $3$-CNFs which are further restricted in certain ways (for example, $3$-CNFs which are required to include a $2$-clause), and 3. identification of appropriate anchors to carry out the induction taking advantage of forced variables (this ingredient in explain in detail in Section 4. In a nutshell, starting with precisely optimal constructions played an essential role in the proving optimality of our constructions.
2 Set Systems
A set system $\mathbb{F}=(V,F)$ is a collection $F$ of sets over a set $V$ of elements. If $V$ is a set of vertices, then $\mathbb{F}$ is interpreted as a graph. If $V$ is as a set of Boolean variables, then $\mathbb{F}$ is interpreted as a monotone CNF. While we regard $\mathbb{F}$ as a set system most of the time, we invite the reader to apply other interpretations to gain additional viewpoints. When we refer to sets in $\mathbb{F}$, we mean the sets in $F$.
@@ 2 Set Systems
We use $n(\mathbb{F})$ to denote the number of elements and $m(\mathbb{F})$ to denote the number of sets of $\mathbb{F}$. We use $k$-set to refer to a set with cardinality $k$. We extend the notation to $k$-clauses and $k$-edges.
We often work with set systems that where the size of the sets is bounded. We say that a set system $\mathbb{F}=(V,F)$ is a $k$-system if each $C\in F$ has cardinality at most $k$. If every set in the set system is a $k$-set, then we call it a $k$-uniform set system.
We now introduce the essential concept of a transversal of a set system.
@@
Let $\mathbb{F}=(V,F)$ be a set system. An $\ell$-transversal of $\mathbb{F}$ is a subset of $V$ of size $\ell$ that intersects every member of $F$. We use transversal to refer to an $\ell$-transversal for some $\ell\geq 0$.
Fact 2.3 (Critical sets of a transversal [13]).
If $\gamma$ is a minimum-size transversal of a set system $\mathbb{F}$, for every $x\in\gamma$ there exists a set $S_{x}$ in $\mathbb{F}$ such that $S_{x}\cap\gamma=\{x\}$. We call $S_{x}$ a critical set for $x$ with respect to $\gamma$.
We parameterize set systems by the minimum size of its transversals to arrive at the notion of $t$-threshold systems.
Definition 2.4 ($t$-threshold set systems).
A set system is called $t$-threshold if all its transversals have size at least $t$.
@@
The transversal number $\tau(\mathbb{H})$ of a set system $\mathbb{H}$ is the cardinality of its minimum-size transversals.
@@
Definition 2.7 (Extremal $t$-threshold $k$-set systems).
We say that a $t$-threshold $k$-system $\mathbb{F}$ is extremal if $\theta_{t}(\mathbb{F})$ is the maximum over all $t$-threshold $k$-systems on $n(\mathbb{F})$ elements.
@@
For $n,k\geq 1$ and $0\leq t\leq n$, we define $\Theta(n,t,k)$ to be equal to $\theta_{t}(\mathbb{H})$ where $\mathbb{H}$ is an extremal $t$-threshold $k$-system over $n$ elements.
The following states that uniform set systems suffice as far as extremal $t$-threshold $k$-set systems are concerned.
@@
For $t\leq n-k+1$ and for every $t$-threshold $k$-system, there is a $t$-threshold $k$-uniform set system which has the same set of $t$-transversals. For t=nk+it=n-k+i and $1\leq i\leq k$, the $t$-threshold (ki+1)(k-i+1)-uniform set system consisting of all (ki+1)(k-i+1)-sets is extremal.
@@ 2.1 Circuits and Transversals
$k$-systems play a key role in the (monotone) circuit complexity of Boolean functions. Our circuits are $\Sigma\Pi\Sigma_{k}$ circuits which are depth-3 circuits where the top gates is an $\mathbf{OR}$ gate and its depth-2 subcircuits are $k$-CNFs. The size of such a circuit is the number of its depth-2 subcircuits. For monotone $\Sigma\Pi\Sigma_{k}$ circuits, the depth-2 subcircuits are $k$-systems.
@@
A Boolean function $f$ on $n$ variables $x_{1},\ldots,x_{n}$ is a $t$-threshold function for $0\leq t\leq n$ provided $f(x_{1},\ldots,x_{n})=1$ if and only if $\Sigma_{i}x_{i}\geq t$.
If a $k$-system $\mathbb{F}$ is a depth-2 subcircuit of a monotone $\Sigma\Pi\Sigma_{k}$ circuit that computes a $t$-threshold function, then $\mathbb{F}$ is a $t$-threshold system. Since no such $\mathbb{F}$ has more than $\Theta(n,t,k)$ $t$-transversals, we conclude the following.
@@
If a monotone $\Sigma\Pi\Sigma_{k}$ circuit computes a $t$-threshold function, then its size is at least $\binom{n}{t}/\Theta(n,t,k)$.
Furthermore, it is known that we only need extremal $t$-threshold $k$-systems as depth-2 subcircuits to construct size-optimal (up to a polynomial factor) monotone $\Sigma\Pi\Sigma_{k}$ circuits for computing $t$-threshold functions. For every $t$-threshold $k$-system $\mathbb{F}$, it is known that there exists a $\Sigma\Pi\Sigma_{k}$ circuit of size ${\mathtt{poly}}(n)\binom{n}{t}/\theta(\mathbb{F})$ for computing a $t$-threshold function. This can be shown by exploiting the symmetries of the threshold functions and monotonicity [4].
@@
There exists a monotone $\Sigma\Pi\Sigma_{k}$ circuit that computes a $t$-threshold function whose size is at most ${\mathtt{poly}}(n)\binom{n}{t}/\Theta(n,t,k)$.
2.2 Connection to Turán Systems
The following definition is adapted from [16, 9].
Definition 2.13 (Turán system and Turán number).
For integers $n\geq\ell\geq k\geq 1$, a $k$-system over an $n$-element set is called an $(n,\ell,k)$-Turán system if every $\ell$-set contains a set of the system. The Turán number ${\Psi}(n,\ell,k)$ is defined as
${\Psi}(n,\ell,k)=\max_{j\leq k}\min m(\mathbb{F}),$
where the minimum is taken over all $(n,\ell,j)$-Turán systems $\mathbb{F}$.
Fact 2.14.
Assume that $2k\leq n-1$. A collection of $k$-sets of an $n$-element set is an $(n,k+1,k)$-Turán system if and only if it is a (nk)(n-k)-threshold $k$-set system.
Proof.
Let $\mathbb{F}$ be an $(n,k+1,k)$-Turán system. Since every $(k+1)$-set includes a $k$-set of $\mathbb{F}$, every (nk1)(n-k-1)-set fails to intersect some member of $\mathbb{F}$. Thus $\mathbb{F}$ is an (nk)(n-k)-threshold $k$-uniform set system.
Conversely, if $\mathbb{F}$ is an (nk)(n-k)-threshold $k$-set system, then no (nk1)(n-k-1)-set is a transversal for $\mathbb{H}$. By considering complements, we conclude $\mathbb{F}$ must then be an $(n,k+1,k)$-Turán system. ∎
Furthermore, we have
Fact 2.15.
$\Theta(n,n-k,k)\;=\;\binom{n}{k}\;-\;{\Psi}(n,k+1,k).$
2.3 Configurations
To prove upper bounds on the number of $t$-transversals of a $t$-threshold system, our arguments will be based on whether certain configurations of sets are present in the set system. In the following, we catalog these configurations.
Definition 2.16 (Triangle configuration).
We say that a set system contains a triangle configuration if it contains three sets of the form $\{a,b,d\}$, $\{b,c,e\}$, and $\{a,c,f\}$ where the set $\{a,b,c\}$ is disjoint from the multiset $\{d,e,f\}$.
Definition 2.17 ( $\mathbf{\exists}$} configuration).
We say that a set system contains a $\mathbf{\exists}$} configuration if it contains four sets of the form $\{a,b,c\}$, $\{a,d,e\}$, $\{b,f,g\}$ and $\{c,h,i\}$ where the set $\{d,e,f,g,h,i\}$ is disjoint from the set $a,b,c$. We say that $\{a,b,c\}$ is the spine and $\{a,d,e\}$, $\{b,f,g\}$ and $\{c,h,i\}$ are the arms of the $\mathbf{\exists}$} configuration.
We sometimes regard a collection of $2$-sets as a graph where each $2$-set is an edge and use graph-theoretic terminology to describe configurations.
Definition 2.18 (Path configuration).
We say that a collections of $2$-sets is a path of length $\ell$ if it contains $\ell$ sets of the form $\{a_{0},a_{1}\}$, $\{a_{1},a_{2}\}$, $\ldots$, $\{a_{\ell-1},a_{\ell}\}$ where $a_{0},a_{1},\ldots,a_{\ell}$ are distinct.
Definition 2.19 (Cycle configuration).
We say that a collection of $2$-sets is a cycle of length $\ell$ if it contains $\ell$ sets of the form $\{a_{0},a_{1}\}$, $\{a_{1},a_{2}\}$, $\ldots$, $\{a_{\ell-1},a_{0}\}$ where $a_{0},a_{1},\ldots,a_{\ell-1}$ are distinct.
3 Constructions for Extremal Threshold System
In this section, we construct $k$-systems over $n\geq 1$ elements for transversal number $0\leq t\leq\lfloor\frac{n}{2}\rfloor$. In the next section, we argue that they are $t$-extremal. We first consider certain constructions for bounded $n$ which will serve as building blocks for general constructions. The basic building blocks come in two flavors: cliques and Turán-style constructions.

### v2

@@ 1 Introduction
What is the largest number of minimum weight satisfying assignments of an $n$-variate monotone 3-CNF with no satisfying assignment of Hamming weight less than $n/2$? We show a tight upper bound of $6^{n/4}$. A construction matching this bound is given by dividing the variables into disjoint blocks of size 4, and for each block including all the monotone 3-clauses. Our proof in fact gives a deterministic algorithm that enumerates these solutions. This problem is closely related to the complexity of Majority function and $k$-SAT.
Circuit complexity.
@@
Paturi, Pudlák, and Zane [13] (hereafter called PPZ) showed that Parity requires $\Sigma^{k}_{3}$ circuits of size $\Omega(2^{n/k})$. In fact, this results in an unrestricted depth-3 lower bound of $\Omega(n^{1/4}2^{\sqrt{n}})$. Both of these lower bounds are known to be tight. Paturi, Pudlák, Saks, and Zane [12] (hereafter called PPSZ) extended the techniques of [13] and showed that recognizing the codewords of a good error-correcting code such as BCH requires $\Sigma^{k}_{3}$ circuits of size $2^{cn/k}$ for a universal constant $c>1$. This remains the best known lower bound for computing any explicit function by $\Sigma^{k}_{3}$ circuits. The core technique in these lower bounds is a combinatorial result about the structure of satisfying assignments of $k$-CNF formulas which can be exploited to solve $k$-SAT.
@@
An algorithm solving $k$-SAT is non-trivial if it has running time $2^{(1-\epsilon_{k})n}$ for some $\epsilon_{k}>0$ which we call savings. PPZ and PPSZ used their combinatorial arguments to give non-trivial $k$-SAT algorithms. PPSZ currently holds the record as the best known $k$-SAT algorithm with savings $c/k$ for a universal constant $c>1$ [14]. The Super Strong Exponential Time Hypothesis (SSETH) states that $\Theta(1/k)$ savings are best possible [17].
@@
The Majority function has long been held as a natural candidate to beat the depth-3 barrier [7]. Unlike Parity, the tightness of natural circuit constructions for Majority is not known. To determine the depth-3 complexity of Majority, Gurumukhani et al. [5] suggested a path forward by considering $\Sigma^{k}_{3}$ circuits. Their central conjecture is that the size of any $\Sigma_{3}^{k}$ circuit computing Majority is $2^{\Omega(n\log(k)/k)}$. Together with a standard random restriction argument to reduce the bottom fan-in, this conjecture implies an unrestricted depth-3 lower bound of $2^{\Omega(\sqrt{n\log n})}$ matching a known upper bound [10]. They pointed out that determining $\Sigma^{k}_{3}$ complexity of Majority is non-trivial even for $k=3$, and proposed this as a first step towards settling the depth-3 complexity of Majority. Results of Lecomte, Ramakrishnan, and Tan [11] can be interpreted as proving the conjecture under the assumption that each depth-2 sub-circuits depends on at most $k$ input bits.
Similar to PPZ and PPSZ, determining the depth-3 complexity is closely tied to $k$-SAT. Gurumukhani et al. [6] showed that proving the conjectured lower bound for $\Sigma^{k}_{3}$ via a particular approach, which they called local enumeration, would break SSETH.
@@
Enum($k$, $t$) is defined as follows. For a natural number $k$ and a parameter $t$, given an $n$-variate $k$-CNF $\mathbb{F}$ with no satisfying assignment of Hamming weight less than $t(n)$, enumerate all satisfying assignments of Hamming weight exactly $t(n)$. Analogously, we can define a restriction of this problem to monotone formulas which we denote by Enum+($k$, $t$).
@@ 1.1 Our contribution
Our main result is an optimal lower bound on the size of any monotone $\Sigma^{3}_{3}$ circuit computing Majority matching the upper bounds conjectured in the literature (see e.g. [4]) which state $\operatorname{Size}^{k}_{3}(\operatorname{\mathsf{Maj}}_{n})\leq 2^{O(n\log(k)/k)}$ and, in particular, $\operatorname{Size}^{3}_{3}(\operatorname{\mathsf{Maj}}_{n})\leq(2/6^{1/4})^{n}\simeq 1.277^{n}$. We prove our lower bound by optimally solving Enum+($3$, $n/2$).
@@
The optimal algorithm for Enum+($3$, $t$) for $t\leq n/2$ is in turn obtained by showing via a constructive argument that extremal monotone $3$-CNF consistent with Majority function (explained below) cannot accept more than $6^{n/4}$ transversals (satisfying assignments) of Hamming weight $\lfloor\frac{n}{2}\rfloor$.
@@
While we do not claim that our techniques are fundamentally new, we believe that our approach to prove matching lower bounds for depth-3 circuits holds promise. The first element of our approach is the top-down technique: if a $\Sigma_{3}^{k}$ circuit computes a Boolean function $f:\{0,1\}^{n}\mapsto\{0,1\}$ and $s$ is the number of its depth-2 circuits (equivalently fan-in of the top OR gate), one of the depth-2 circuits (which is a $k$-CNF) must accept at least $|f^{-1}(1)|/s$ members of $f^{-1}(1)$. This brings our focus to the study of extremal $k$-CNFs with respect to $f$. A $k$-CNF $\mathbb{F}$ is extremal with respect to $f$ if ${\textsc{Sat}}(\mathbb{F})\subseteq f^{-1}(1)$ ($\mathbb{F}$ is consistent with $f$) and $|{\textsc{Sat}}(\mathbb{F})|$ is the maximum among such $k$-CNF. Such extremal $k$-CNFs can also used for constructing $\Sigma_{3}^{k}$ circuits (almost) optimally depending on the symmetries of $f$. Thus, the study of $k$-CNFs which are extremal with respect to a Boolean function is of fundamental concern.
@@
What do we know of $k$-CNFs which are extremal with respect to Majority function or more generally with respect to Boolean threshold functions? Our intuition (based on optimal constructions for small $n$) suggested that a modularity principle might hold for extremal $k$-CNF with respect to Majority. We say that the modularity principle holds for a Boolean function if one of its extremal $k$-CNF $\mathbb{F}$ can be expressed as the conjunction of $k$-CNFs $\mathbb{F}^{\prime}_{i}$ where $\mathbb{F}^{\prime}_{i}$ is supported on a bounded number of variables and the supports of $\mathbb{F}^{\prime}_{i}$ are disjoint. Although we lacked a formal method to determine if the Majority function is modular a priori, we used the modularity principle to guide our search for extremal 3-CNFs of Boolean threshold functions for small values of $n$ and then generalized the constructions for all $n$.
Our final step is to apply induction to prove the optimality of our constructions. Several ingredients helped the induction to succeed:
  1. 1.
    Inductive structure in the constructions of extremal monotone $3$-CNFs for general and restricted $3$-CNFs (for example, $3$-CNFs which are required to include a $2$-clause),
  2. 2.
    Correct guesses of extremal (restricted and unrestricted) $3$-CNFs for Boolean threshold functions with threshold less or equal to $n/2$, and
  3. 3.
    Intricate division of formulas (which are consistent with threshold functions) into subclasses.
In a nutshell, we used an approach of upper bounds guiding the lower bounds where guessing the optimal constructions played an essential role in the proving their optimality. Our approach to the proof leads to an algorithm to enumerate the transversals.
2 Monotone CNFs and Transversals
A CNF (for Conjunctive Normal Form) formula $\mathbb{F}=(V,F)$ is a collection $F$ of clauses (where each clause is a set of literals) over a set $V$ of Boolean variables. We assume without loss of generality that the literals in the clause correspond to distinct variables.
@@ 2 Monotone CNFs and Transversals
We use $n(\mathbb{F})$ to denote the number of variables and $m(\mathbb{F})$ to denote the number of clauses of $\mathbb{F}$. We use $k$-clause to refer to a clause with cardinality $k$. We extend the notation to $k$-sets and $k$-edges.
We often work with CNFs that where the size of the clauses is bounded. We say that a CNF $\mathbb{F}=(V,F)$ is a $k$-CNF if each $C\in F$ has cardinality at most $k$. If every clause in the CNF is a $k$-clause, then we call it a $k$-uniform CNF.
In this paper, we will only deal with monotone CNFs, where each clause consists of only positive literals. A clause which only contains positive literals is called a monotone clause. If $V$ is a set of vertices, then a monotone CNF $\mathbb{F}$ is interpreted as a hypergraph. If $V$ is a set of elements, then $\mathbb{F}$ is interpreted as a set system. While we speak in the language of CNFs $\mathbb{F}$, we invite the reader to apply other interpretations to gain additional viewpoints. When we refer to clauses in $\mathbb{F}$, we mean the clauses in $F$.
We now introduce the concept of a transversal (or satisfying assignment) of a monotone CNF.
@@
Let $\mathbb{F}=(V,F)$ be a monotone CNF. An $\ell$-transversal of $\mathbb{F}$ is a subset of $V$ of size $\ell$ that intersects every member of $F$. We use transversal to refer to an $\ell$-transversal for some $\ell\geq 0$.
Fact 2.3 (Critical clauses of a transversal [13]).
If $\gamma$ is a minimum-size transversal of a monotone CNF $\mathbb{F}$, for every $x\in\gamma$ there exists a clause $C_{x}$ in $\mathbb{F}$ such that $C_{x}\cap\gamma=\{x\}$. We call $C_{x}$ a critical clause for $x$ with respect to $\gamma$.
We parameterize monotone CNFs by the minimum size of its transversals to arrive at the notion of $t$-threshold systems.
Definition 2.4 ($t$-threshold CNFs).
A monotone CNF is called a $t$-threshold CNF if all its transversals have size at least $t$.
@@
The transversal number $\tau(\mathbb{F})$ of a monotone CNF $\mathbb{F}$ is the cardinality of its minimum-size transversals.
@@
Definition 2.7 (Extremal $t$-threshold $k$-CNFs).
We say that a $t$-threshold monotone $k$-CNF $\mathbb{F}$ is extremal if $\theta_{t}(\mathbb{F})$ is the maximum over all $t$-threshold monotone $k$-CNFs over $n(\mathbb{F})$ variables.
@@
For $n,k\geq 1$ and $0\leq t\leq n$, we define $\Theta(n,t,k)$ to be equal to $\theta_{t}(\mathbb{H})$ where $\mathbb{H}$ is an extremal $t$-threshold monotone $k$-CNF over $n$ variables.
Note that uniform $k$-CNFs suffice as far as extremal $t$-threshold monotone $k$-CNFs are concerned.
@@
For $t\leq n-k+1$ and for every $t$-threshold monotone $k$-CNF, there is a $t$-threshold monotone $k$-uniform CNF which has the same set of $t$-transversals. For t=nk+it=n-k+i and $1\leq i\leq k$, the $t$-threshold monotone (ki+1)(k-i+1)-uniform CNF consisting of all (ki+1)(k-i+1)-sets is extremal.
@@ 2.1 Circuits and Transversals
Monotone $k$-CNFs play a key role in the monotone circuit complexity of Boolean functions. We consider $\Sigma_{3}^{k}$ circuits which are depth-3 circuits where the top gate is an $\mathbf{OR}$ gate and its depth-2 sub-circuits are $k$-CNFs. The size of such a circuit is the number of its depth-2 sub-circuits. For monotone $\Sigma_{3}^{k}$ circuits, the depth-2 sub-circuits are monotone $k$-CNFs.
@@
On $n$ variables $x_{1},\ldots,x_{n}$, a $t$-threshold function $f$ for $0\leq t\leq n$ is a Boolean function with $f(x_{1},\ldots,x_{n})=1$ if and only if $\Sigma_{i}x_{i}\geq t$.
If a monotone $k$-CNF $\mathbb{F}$ is a depth-2 sub-circuit of a monotone $\Sigma^{k}_{3}$ circuit that computes a $t$-threshold function, then $\mathbb{F}$ is a $t$-threshold CNF. Since no such $\mathbb{F}$ has more than $\Theta(n,t,k)$ $t$-transversals, we conclude the following.
@@
If a monotone $\Sigma^{k}_{3}$ circuit computes a $t$-threshold function, then its size is at least $\binom{n}{t}/\Theta(n,t,k)$.
Furthermore, it is known that we only need extremal $t$-threshold monotone $k$-CNFs as depth-2 sub-circuits to construct size-optimal (up to a polynomial factor) monotone $\Sigma^{k}_{3}$ circuits for computing $t$-threshold functions. For every $t$-threshold monotone $k$-CNF $\mathbb{F}$, it is known that there exists a $\Sigma^{k}_{3}$ circuit of size ${\mathtt{poly}}(n)\binom{n}{t}/\theta(\mathbb{F})$ for computing a $t$-threshold function. This can be shown by exploiting the symmetries of the threshold functions and monotonicity [4].
@@
There exists a monotone $\Sigma^{k}_{3}$ circuit that computes a $t$-threshold function whose size is at most ${\mathtt{poly}}(n)\binom{n}{t}/\Theta(n,t,k)$.
3 Constructions for Extremal Threshold CNFs
In this section, we construct $k$-CNFs over $n\geq 1$ variables for transversal number $0\leq t\leq\lfloor\frac{n}{2}\rfloor$. In the next section, we argue that they are $t$-extremal. We first consider certain constructions for bounded $n$ which will serve as building blocks for general constructions. The basic building blocks come in two flavors: cliques and Turán-style constructions.